OpenStudy (anonymous):

Need help Graph the equation by plotting the points? y=1/3x-2 (3, ) (0, ) 2nd ? Find the slope of the following line? 4y=x+8

OpenStudy (anonymous):

first of all you can draw your equation at this website .. most of users here do that .. http://www.wolframalpha.com second you just have to sub x in your equation and get y values (3,1/7) , (0,-0.5) third thing .. and its most important y= mx +b this is the standard form of first degree equation with two variables where m = slope ,b= y intercept 4y=x+8 y=(1/4)x+2 slope = (1/4) you are welcome